OpenStudy (anonymous):

Complex number problem If z= 7-24i, a<= lz+2l <=b , what is the maximum value of b-a lz+2l denotes the absolute value of z+2

OpenStudy (michele_laino):

we have: \(z+2=9-24i\) so:\[\left| {z + 2} \right| = \sqrt {81 + 576} = 25.63...\] therefore, we can write: \[24 \leqslant \left| {z + 2} \right| \leqslant 26\]

OpenStudy (michele_laino):

better is: \[25 \leqslant \left| {z + 2} \right| \leqslant 26\]

OpenStudy (anonymous):

So what is the maximum value? I tried 1 and 2 but it isn't the correct answer.

OpenStudy (anonymous):

@Michele_Laino

geerky42 (geerky42):

Maximum? You mean minimum? Because for me, answer would be something like \(\infty\).

geerky42 (geerky42):

Do you mean \(a-b\)? can you recheck question?